OpenStudy (anonymous):

Hey !:) how would you find the integral of a function with no specific u value..such as...?

OpenStudy (anonymous):

\[\int\limits_{?}^{?} x ^{3}-6x-20divx+5\]

OpenStudy (anonymous):

\[∫??x3−6x−20 \div x+5\]

OpenStudy (anonymous):

Just rewrite it as: \[\int\limits x^3-6x-20x^{-1}+5 dx\] Integrate term by term.

OpenStudy (anonymous):

can u show more steps..i always get stcuk with integration by parts...:(

OpenStudy (anonymous):

so it wud b..x^3/3-6x^2/2+5x +c

OpenStudy (anonymous):

Well if you have 20/x then it would be 20ln|x| because you know that the derivative of ln(x)=1/x so the integral of 1/x is ln|x| xP

OpenStudy (anonymous):

But the rest is right :)

OpenStudy (anonymous):

so the deiv rules apply to integral rules too??

OpenStudy (anonymous):

They work in reverse so if you have the derivative of x is 1. But the integral of 1 is x. They are inverse operations.

OpenStudy (anonymous):

ohh right...thanks a lotttt